# A comprehensive survey on deep active learning and its applications in medical image analysis

### v1

@@ 3.1.1 Multiple Inferences
To mitigate over-confidence, a common strategy is to run the model multiple times with some perturbations and calculate the classic uncertainty metrics with the mean probability. The main idea is to reduce the bias introduced by network architectures or training data. These biases often contribute to the over-confidence issue. Three methods can be employed for multiple inferences: Monte Carlo dropout (MC dropout), model ensemble, and data augmentation. The first two perturb the model parameters, and the last perturb the input data.
@@
MC dropout randomly discards some neurons in the deep model during each inference (Gal and Ghahramani, 2016). With MC dropout enabled, the model runs multiple times to get different predictions. Gal et al. (2017) was the pioneering work of deep AL. They were also the first to use MC dropout in computing uncertainty metrics like entropy, standard deviation, and Bayesian active learning disagreement (BALD) (Houlsby et al., 2011). Results showed that MC Dropout could significantly improve the performance of uncertainty-based deep AL.
@@
Model ensemble trains multiple models to get numerous predictions during inference. Beluch et al. (2018) conducted a detailed comparison of models ensemble and MC dropout in uncertainty-based AL. Results demonstrated that the model ensemble performs better. However, model ensemble requires significant training overhead in DL. To reduce the computational costs, snapshot ensemble (Huang et al., 2017) obtained multiple models in a single run with cyclic learning rate decay. An early attempt in Beluch et al. (2018) showed that snapshot ensemble leads to worse performance than model ensemble. Jung et al. (2023) improved the snapshot ensemble by maintaining the same optimization trajectory in different AL rounds, along with parameter regularization. Results showed that the improved snapshot ensemble outperforms the model ensemble. Additionally, Nath et al. (2021) employed stein variational gradient descent to train an ensemble of models, aiming to ensure diversity among them.
Data augmentation produces different versions of input data with random transformations. Then, multiple predictions were obtained by running the model with different augmentations during inference. In point cloud semantic segmentation, Hu et al. (2022) randomly augmented the input point clouds multiple times, then calculated the entropy of average prediction probability for each point after registration and correspondence estimation. Liu et al. (2022a) also applied random augmentations to point clouds, but they used variance as the uncertainty metric.
3.1.2 Disagreement-based Uncertainty
Disagreements between different inferences of the same sample could also be a measure of uncertainty. Samples with higher disagreement indicate higher uncertainty and are suitable for annotation. The principle of methods in this section aligns with the previous section, which aims to mitigate over-confidence by introducing perturbations during multiple inferences. However, methods in the previous section focus on improving the uncertainty estimation of classical metrics, while methods in this section leverage the disagreements between different prediction results. In this line of research, we can build disagreement from the perspective of model and data.
@@ 3.1.2 Disagreement-based Uncertainty
Model disagreement: We can utilize the disagreement between the outputs of different models, also known as Query-by-Committee (QBC) (Seung et al., 1992). Suggestive annotation (SA) (Yang et al., 2017) trained multiple segmentation networks with bootstrapping. The variance among these models is used as the disagreement metric. Mackowiak et al. (2018) adopted the vote entropy between different MC dropout inferences as the disagreement metric. Peng et al. (2021) trained teacher and student models through knowledge distillation and used the L2 distance of segmentation predictions as the disagreement metric. In polyp segmentation of capsule colonoscopy, Bai et al. (2022) trained multiple decoders using class activation maps (CAMs) (Zhou et al., 2016) generated by a classification network. They further proposed model disagreement and CAM disagreement for sample selection. Model disagreement included entropy of prediction probabilities and Dice between outputs of different decoders, while CAM disagreement measured the Dice between CAMs and outputs of all decoders. This method selected samples with high model disagreement and CAM disagreement for annotation. However, samples with low model disagreement but high CAM disagreement were treated as pseudo-labels for semi-supervised training. In rib fracture detection, Huang et al. (2020) adopted Hausdorff distance to measure the disagreements between different CAMs.
@@
Data disagreement: Since training multiple models can be computationally expensive, measuring the disagreements between different perturbations of input data is also helpful in AL. Kullback-Leibler (KL) divergence is a commonly used metric for quantifying disagreement. Wu et al. (2021b) computed KL divergence between different versions of augmentations as the disagreement measure. Siddiqui et al. (2020) measured the disagreement with KL divergence between predictions of different viewpoints in 3D scenes. In point cloud segmentation, Hu et al. (2022) employed KL divergence to measure the disagreement between predictions of different frames. Additionally, recent works have adopted alternative metrics to calculate disagreement. Lyu et al. (2023) proposed input-end committee, which constructs different predictions by randomly augmenting the input data. They further measured the classification and localization disagreements between different predictions with cross-entropy and variance, respectively. Parvaneh et al. (2022) interpolated the unlabeled samples and labeled prototypes in the feature space. If the prediction of the interpolated sample disagrees with the label of the prototype, it indicates that the unlabeled samples may introduce new features. Then, these unlabeled samples should be sent for annotation. Results showed advancements across various datasets and settings. Besides, some works explored the prediction disagreement within a local area inside an image. In object detection, Aghdam et al. (2019) assumed that the disagreement between probabilities in the neighborhood of wrongly predicted pixels should be high. They adopted BALD as the disagreement metric.
@@ 3.1.3 Uncertainty-aware Model
The main idea of uncertainty-aware models is to transform commonly used deterministic models into probabilistic models. In this way, the network no longer outputs a single point estimate but instead provides a distribution of possible predictions, thus mitigating over-confidence. This approach only requires a single pass of the deep model, significantly reducing computational and time costs during inference. Related works mainly fall into two categories: evidential deep learning (EDL) and mixture density networks (MDN).
Evidential deep learning replaces the Softmax distribution with a Dirichlet distribution (Sensoy et al., 2018). The network’s output is interpreted as the parameters of a Dirichlet distribution, so the predictions followed the Dirichlet distribution. The Dirichlet distribution will be sharp if the model is confident about the predictions. Otherwise, it will be flat. To make EDL compatible with object detection tasks, Park et al. (2023) introduced a model evidence head to scale the parameters of the Dirichlet distribution adaptively, which enhanced training stability. They first calculated the epistemic uncertainty for each detection box. Then, the sample-level uncertainty was obtained through hierarchical uncertainty aggregation. Sun et al. (2023) also applied EDL in scene graph generation, achieving near full-supervision performance with only about 10% of the labeling cost.
@@
3.1.4 Gradient-based Uncertainty
Gradient-based optimization is essential for deep learning. The gradient of each sample reflects its contribution to the change of model parameters. A larger gradient length indicates a tremendous change of parameters by the sample, thus implying high uncertainty. Furthermore, gradients are independent of predictive probabilities, making them less susceptible to over-confidence. In deep active learning, three metrics are frequently used as gradient-based uncertainty: gradients, Fisher information (FI), and influence functions.
Gradient: A larger gradient norm (i.e., gradient length) denotes a greater influence on model parameters, indicating higher uncertainty in AL. Ash et al. (2020) proposed batch active learning by diverse gradient embeddings (BADGE). They calculated the gradients only for the parameters of the network’s final layer, with the most confident classes as pseudo labels in gradient computation. Then, k-Means++ is performed on gradient embeddings for sample selection. Results showed competitive performances of BADGE across diverse datasets, network architectures, and hyperparameter settings. Wang et al. (2022b) proved that a larger gradient norm corresponds to a lower upper bound of test loss. Thus, they employed expected empirical loss and entropy loss for gradient computation, which both obviate the necessity for labels. The former was a sum of the losses for each class, which was weighted by the corresponding probability. The latter was the entropy of probabilities of all classes. In MRI brain tumor segmentation, Dai et al. (2020) employed gradients for active learning. They first trained a variational autoencoder (VAE) (Kingma and Welling, 2013) to learn the data manifold. Then, they trained a segmentation model and calculated gradients of Dice loss using available labeled data. The sample selection was guided by the gradient projected onto the data manifold. Their extended work (Dai et al., 2022) further demonstrated superior performance in MRI whole brain segmentation.
Fisher information: As the expectation of the gradient’s covariance matrix, FI reflects overall uncertainty according to the data distribution to model parameters. Annotating samples with higher FI helps the model converge faster toward optimal parameters. FI has already succeeded in AL of machine learning models (Chaudhuri et al., 2015; Sourati et al., 2017). However, the computation cost of FI grows quadratically with the increase of model parameters, which is unacceptable for deep active learning. Sourati et al. (2018) and their extended work (Sourati et al., 2019) were the first to incorporate FI into deep active learning. They used the average gradients of each layer to calculate the FI matrix, thus reducing the computation cost. This method outperformed competitors in brain extraction across different age groups and pathological conditions. Additionally, Ash et al. (2021) only computed the FI matrix for the network’s last layer.
Influence functions: Liu et al. (2021b) employed influence functions (Koh and Liang, 2017) to select samples that bring the most positive impact on model performance. The method used expected empirical loss to calculate the gradient since influence functions also require gradient computations.
@@ 3.1.5 Adversarial-based Uncertainty
Adversarial samples are created by adding carefully designed perturbations to normal samples by attacking the deep models (Goodfellow et al., 2014b). The differences between adversarial and original samples are nearly indiscernible to the human eye. However, deep models would produce extremely confident wrong predictions for adversarial samples. The reason is that adversarial attacks push the original samples to the other side of the decision boundary with minimal cost, resulting in visually negligible changes but significantly different predictions. From this perspective, the strength of adversarial attacks reflects the sample’s distance to the decision boundary (Heo et al., 2019). A small perturbation indicates that the sample is closer to the decision boundary and, thus, is considered more uncertain. Ducoffe and Precioso (2018) adopted the DeepFool algorithm (Moosavi-Dezfooli et al., 2016) for adversarial attacks. Samples with small adversarial perturbations are requested for labeling. Rangwani et al. (2021) attacked the deep model by maximizing the KL divergence between predictions of adversarial and original samples while the strength of perturbation is limited.
Adversarial training involves alternating training between feature extractors and multiple classifiers. The objectives of training feature extractors and classifiers are conflicting. Multi-round adversarial training increases the disagreements between classifiers, uncovering uncertain samples hidden by over-confidence. In object detection, Yuan et al. (2021) and their extended work (Wan et al., 2023) used two classifiers for adversarial training on both labeled and unlabeled datasets. The first step is to fix the feature extractor and tune the two classifiers. The more classifiers disagreed the more uncertain samples were exposed. Then, they fixed the classifiers and tuned the feature extractor with opposite objectives, aiming to narrow the distribution gap between labeled and unlabeled samples. After multiple rounds of alternating training, samples with the highest disagreements between classifiers are sent for annotation. Fu et al. (2021) adversarially trained multiple classifiers to maximize their disagreement. The standard deviation between different predictions was considered the uncertainty metric.
3.1.6 Performance Estimation
In this section, the uncertainty metrics are direct performance estimations of the current task. There are two types of such metrics: test loss or task-specific evaluation metrics. These metrics reflect the level of prediction error. For instance, a low Dice score suggests the model failed to produce accurate segmentation. Request annotations for these samples would be beneficial for improving the model’s performance. However, instead of calculating these metrics precisely, we can only estimate them without the ground truths. There are primarily two methods for estimating performance: surrogate metrics and learnable performance estimation.
Surrogate metrics are widely used in this line of research. For example, these metrics may be upper or lower bounds for loss or some evaluation metrics. Huang et al. (2021) found that within limited training iterations, the loss of a sample is bounded by the norm of the difference between the initial and final network outputs. Inspired by this, they proposed cyclic output discrepancy (COD) as the difference in model output between two consecutive annotation rounds. Results indicated that a higher COD is associated with higher loss. Therefore, they opted for samples with high COD. They may also demonstrate a linear correlation with the evaluation metrics with post-hoc validation. Shen et al. (2020) calculated the intersection over union (IoU) of all predictions by MC dropout. They found a strong linear correlation between this IoU and the real Dice coefficient. Zhao et al. (2021) calculated the average Dice coefficient between the predictions of the intermediate layers and final layer through deep supervision. They also found a linear correlation between this average Dice and the real Dice coefficient. Results showed competitive performance in skin lesion segmentation and X-ray hand bone segmentation.
Learnable performance estimation: Additionally, we can train auxiliary neural network modules to predict the performance metrics. As one of the most representative works in this line of research, learning loss for active learning (LL4AL) (Yoo and Kweon, 2019) trained an additional module to predict the loss value of a sample without its label. Since loss indicates the quality of network predictions, the predicted loss is a natural uncertainty metric for sample selection. Results showed that predicted and actual losses are strongly correlated. LL4AL also outperformed several AL baselines. In lung nodule detection with CT scans, Liu et al. (2020) built upon LL4AL to predict the loss of each sample and bounding box. In diagnosing COVID-19, Wu et al. (2021b) adopted both the predicted loss and the disagreements between different predictions for sample selection. Since AL focuses only on uncertainty ranking of the unlabeled samples, Kim et al. (2021) relaxed the loss regression to loss ranking prediction. Thus, they replaced the loss regressor in LL4AL with the ranker in RankCGAN (Saquil et al., 2018). Results showed that loss ranking prediction outperforms LL4AL.

### v3

@@ 3.1.1 Uncertainty with Multiple Inferences
To mitigate over-confidence, a common strategy for uncertainty-based AL is to run the model multiple times under perturbations. The main idea is to reduce the bias introduced by network architectures or training data. These biases often contribute to the over-confidence issue. Two approaches are often used to utilize multiple inference results for AL. The first is to calculate the classic uncertainty metrics with the average probability of multiple inferences. Averaging the prediction probabilities of multiple inferences helps to reduce individual bias that causes over-confidence. The other approach takes the disagreement between different prediction results as uncertainty quantification. Samples with higher disagreement indicate higher uncertainty and are suitable for annotation in AL.
In this section, we will introduce four types of methods for AL with multiple inferences: Monte Carlo dropout (MC dropout), model ensemble, model disagreement, and data disagreement. The first two used the average probability of results of multiple inferences to calculate the uncertainty metrics, like entropy and margin. The last two are based on disagreement. For the source of perturbation, the first three perturb the model parameters while the last perturb the input data.
@@
MC dropout randomly discards certain neurons in the deep model during each inference (Gal and Ghahramani, 2016). With MC dropout enabled, the model runs multiple times to get different predictions. Gal et al. (2017) was the pioneering work of deep AL. They were the first to use MC dropout in computing uncertainty metrics like entropy, standard deviation, and Bayesian active learning disagreement (BALD) (Houlsby et al., 2011). Results showed that MC Dropout could significantly improve the performance of uncertainty-based deep AL. Besides, they were also among the first to apply deep AL in medical image analysis. In the skin lesion analysis dataset ISIC 2016, they found that BALD consistently outperformed the random baseline. In brain cell type classification, Yuan et al. (2020b) calculated entropy using the average probability of multiple MC dropout runs. Gu et al. (2018) adopted the variance of multiple MC dropout runs as the uncertainty metric in the classification of confocal endomicroscopy and gastrointestinal endoscopy.
@@
Model ensemble trains multiple models to get numerous predictions during inference. Beluch et al. (2018) conducted a detailed comparison of models ensemble and MC dropout in uncertainty-based AL. Results in the standard datasets demonstrated that the model ensemble performs better. For AL in diagnosing diabetic retinopathy, the proposed method achieved significant improvement compared to the random baseline. However, model ensemble requires significant training overhead in DL. To reduce the computational costs, snapshot ensemble (Huang et al., 2017) obtained multiple models in a single run with cyclic learning rate decay. An early attempt in Beluch et al. (2018) showed that snapshot ensemble leads to worse performance than model ensemble. Jung et al. (2023) improved the snapshot ensemble by maintaining the same optimization trajectory in different AL rounds, along with parameter regularization. Results showed that the improved snapshot ensemble outperforms the model ensemble. Nath et al. (2021) employed stein variational gradient descent to train an ensemble of models, aiming to ensure diversity. Their proposed method showed advantages to other competitors in segmenting the pancreas and tumor on CT and hippocampus on MRI.
@@
Model disagreement: We can utilize the disagreement between the outputs of different models, which can be also referred to as Query-by-Committee (QBC) (Seung et al., 1992). This type of method was widely used in AL for medical image analysis. Suggestive annotation (SA) is the pioneering work of AL for medical image analysis (Yang et al., 2017). They trained multiple segmentation networks with bootstrapping. Variance among these models is used as the disagreement metric. SA demonstrated superior performance in segmenting glands on pathological images and lymph nodes on ultrasound images. In abdominal multi-organ segmentation, Qu et al. (2023a) trained three different segmentation models and adopted the variance between their predictions. In carotid intima-media segmentation for ultrasound images, Tang et al. (2023) selected samples with the highest Kullback-Leibler (KL) divergence between the predictions of teacher and student models for annotation. In polyp segmentation of capsule colonoscopy, Bai et al. (2022) trained multiple decoders using class activation maps (CAMs) (Zhou et al., 2016) generated by a classification network. They further proposed model disagreement and CAM disagreement for sample selection. Model disagreement included entropy of prediction probabilities and Dice between outputs of different decoders, while CAM disagreement measured the Dice between CAMs and outputs of all decoders. This method selected samples with high model disagreement and CAM disagreement for annotation. However, samples with low model disagreement but high CAM disagreement were treated as pseudo-labels for semi-supervised training. In rib fracture detection, Huang et al. (2020) adopted Hausdorff distance to measure the disagreements between different CAMs. Besides, Mackowiak et al. (2018) adopted vote entropy between different MC dropout inferences as the disagreement metric.
@@
Data disagreement: Since training multiple models can be computationally expensive, measuring the disagreements between different perturbations of input data is also helpful in AL. Kullback-Leibler (KL) divergence is a commonly used metric for quantifying disagreement. In COVID diagnosis, Wu et al. (2021) computed KL divergence between different versions of augmentations as the disagreement measure to select informative CT scans for annotation. Siddiqui et al. (2020) calculated the KL divergence between predictions of different viewpoints in 3D scenes to select informative regions for AL. Additionally, recent works have adopted alternative metrics to calculate disagreement. Lyu et al. (2023) proposed input-end committee, which randomly augmented the input data to get multiple predictions. They further measured the classification and localization disagreements between different predictions with cross-entropy and variance, respectively. Parvaneh et al. (2022) interpolated the unlabeled samples and labeled prototypes in the feature space. If the interpolated sample’s prediction disagrees with the corresponding prototype’s label, it indicates that the unlabeled samples introduce new features. Thus, these unlabeled samples should be sent for annotation. Results showed advancements across various datasets and settings.
3.1.2 Gradient-based Uncertainty
Gradient-based optimization is the cornerstone of DL-based medical image analysis. The gradient of each sample reflects its contribution to the change of model parameters. A larger gradient length indicates a tremendous change of parameters by the sample, thus implying high uncertainty. Furthermore, gradients are independent of predictive probabilities, which makes them less susceptible to over-confidence. Three metrics that are frequently used as gradient-based uncertainty: gradients, Fisher information, and influence functions. It should be noted that the gradient computation in this section did not use the ground truth labels which are unavailable for unlabeled samples. Instead, the corresponding methods either used supervised loss with pseudo-labels (e.g. cross-entropy loss with pseudo-labels) or unsupervised loss (e.g. entropy loss), thereby making the gradient computation independent of the true labels.
Gradient: A larger gradient norm (i.e., gradient length) denotes a greater influence on model parameters, indicating higher uncertainty in AL. As an early attempt, Otálora et al. (2017) adopted the classic expected gradient length (Settles et al., 2007) to select valuable samples for annotation in exudate classification of eye fundus images. As a popular and pioneering work in the DL era, Ash et al. (2020) proposed batch active learning by diverse gradient embeddings (BADGE). They calculated the gradients only for the parameters of the network’s final layer, with the most confident classes as pseudo labels in gradient computation. Then, k-Means++ is performed on gradient embeddings for sample selection. Results showed competitive performances of BADGE across diverse datasets, network architectures, and hyperparameter settings. Gradient has been widely used in active learning of medical image analysis. Aklilu and Yeung (2022) extended the BADGE framework into semantic segmentation of laparoscopic surgical images. Wang et al. (2022b) proved mathematically that a larger gradient norm corresponds to a lower upper bound of test loss. Thus, they employed expected empirical loss and entropy loss for gradient computation, which both obviate the necessity for labels. The former is the weighted sum of the losses of each class and the class probabilities, which are as follows:
$\mathcal{L}_{exp}(x)=\sum\limits_{i=1}^{C}\left[p_{i}\cdot\mathcal{L}(x,y_{i};\theta)\right]$ (5)
where $y_{i}$ is the label of class $i$. The entropy loss is solely based on the probabilities of all classes, which are as follows:
$\mathcal{L}_{ent}(x)=-\sum\limits_{i=1}^{C}p_{i}\log p_{i}$ (6)
The proposed method outperformed other comparative methods in cryo-electron tomography (cyro-ET) subtomogram classification. Besides, Dai et al. (2020) proposed a new gradients-based active learning method in MRI brain tumor segmentation. They first trained a variational autoencoder (VAE) (Kingma and Welling, 2013) to learn the data manifold. Then, they trained a segmentation model and calculated gradients of Dice loss using available labeled data. The sample selection was guided by the gradient projected onto the data manifold. Their extended work (Dai et al., 2022) further demonstrated superior performance in MRI whole brain segmentation.
Fisher information is effective in AL of machine learning models (Chaudhuri et al., 2015; Sourati et al., 2017). Fisher information (FI) reflects the overall uncertainty of model parameters according to data distribution. FI is defined as the expectation of the squared gradients with respect to the model parameter, the formulation is as follows:
$\mathcal{I}_{Fisher}(x;\theta)=\mathbb{E}_{y}\left[\nabla_{\theta}^{2}\mathcal{L}(x,y;\theta)\right]$ (7)
where $\mathcal{I}$ is the notation of Fisher information. The trace of the inverse of FI often serves as the objective for AL:
$\underset{D^{q}\subset D^{u}}{\arg\min}\operatorname{Tr}\left[\left(\sum_{x\in D^{q}}\mathcal{I}_{Fisher}(x;\theta)\right)^{-1}\left(\sum_{x\in D^{u}}\mathcal{I}_{Fisher}(x;\theta)\right)\right]$ (8)
By solving Eq. 8, the selected samples could help the model converge faster toward optimal parameters. However, the computation cost of FI-based methods grows quadratically with the increase of model parameters, which is unacceptable for deep active learning. Sourati et al. (2018) and their extended work (Sourati et al., 2019) were the first to incorporate FI into deep active learning of medical image analysis. They used the average gradients of each layer to calculate the FI matrix, thus reducing the computation cost. Due to the absence of ground truth labels, they adopt the expected empirical loss (i.e., Eq. 5) for gradient computation. This method outperformed competitors in brain extraction across different age groups and pathological conditions. Additionally, Ash et al. (2021) only computed the FI matrix for the network’s last layer. The gradient computation of this work is the same as that of BADGE (Ash et al., 2020).
Influence function: Liu et al. (2021) employed influence function (Koh and Liang, 2017) to select samples that bring the most positive impact on model performance. The influence function of an unlabeled sample is defined as follows (Weisberg and Cook, 1982):
$\mathcal{I}_{Influence}(x;D^{l})=-\left(\sum_{(x,y)\in D^{l}}\nabla_{\theta}\mathcal{L}(x,y;\theta)\right)H_{\theta}^{-1}\nabla_{\theta}\mathcal{L}(x,y;\theta)$ (9)
where $H_{\theta}^{-1}$ is the Hessian matrix of the labeled set, $H_{\theta}=\sum_{(x,y)\in D^{l}}\nabla_{\theta}^{2}\mathcal{L}(x,y;\theta)$. In Eq. 9, the gradients of the first (i.e., the sum of gradients of labeled samples) and the second term (i.e., the Hessian matrix) can be derived with the ground truth label. For the third term, they replaced the true gradients with the gradients of the expected empirical loss due to the unavailability of the ground truth label.
3.1.3 Performance Estimation
In this section, uncertainty metrics are estimations of the current task’s performance. There are two types of such metrics: test loss or task-specific evaluation metrics. These metrics reflect the level of prediction error. For instance, a low Dice score in tumor segmentation for a patient suggests the model failed to produce accurate segmentation. Request annotations for these samples would be beneficial for improving the model’s performance. However, due to the unavailability of ground truth labels, we can only estimate these metrics instead of calculating them precisely. There are primarily two methods for estimating performance: surrogate metrics and learnable performance estimation.
Surrogate metrics are widely used in active learning of medical image analysis. For example, these metrics could be upper or lower bounds for loss or task-specific evaluation metrics. In breast cancer segmentation on immunohistochemistry images, Shen et al. (2020) calculated the intersection over union (IoU) of all predictions by MC dropout. They found a strong linear correlation between this IoU and the real Dice coefficient. In skin lesion and X-ray hand bone segmentation, Zhao et al. (2021) calculated the average Dice coefficient between the predictions of the intermediate layers and final layer through deep supervision. They found a linear correlation between this average Dice and the real Dice coefficient. Besides, Huang et al. (2021) found that within limited training iterations, the loss of a sample is bounded by the norm of the difference between the initial and final network outputs. Inspired by this, they proposed cyclic output discrepancy (COD) as the difference in model output between two consecutive annotation rounds. Results indicated that a higher COD is associated with higher loss. Therefore, they opted for samples with high COD. They also demonstrate a linear correlation with the evaluation metrics with post-hoc validation.
Learnable performance estimation: We can train auxiliary neural network modules to predict the performance metrics. As one of the most representative works in this line of research, learning loss for active learning (LLAL) (Yoo and Kweon, 2019) trained an additional module to predict the loss value of a sample without its label. Since loss indicates the quality of network predictions, the predicted loss is a natural uncertainty metric for sample selection. Results showed that predicted and actual losses are strongly correlated. The proposed method also outperformed several AL baselines. In lung nodule detection with CT scans, Liu et al. (2020) built upon LLAL to predict the loss of each sample and bounding box. In COVID diagnosis, Wu et al. (2021) adopted both the predicted loss and the disagreements between different predictions for sample selection. Wu et al. (2022c) further combined the loss prediction and sample diversity in the federated active learning of COVID diagnosis and colonoscopy polyp analysis. Since AL focuses only on uncertainty ranking of the unlabeled samples, Kim et al. (2021) relaxed the loss regression to loss ranking prediction. Thus, they replaced the loss regressor in LLAL with the ranker in RankCGAN (Saquil et al., 2018). Results showed that loss ranking prediction outperforms the actual loss regression in LLAL. Zhou et al. (2021b) and their subsequent work (Zhou et al., 2022) introduced a quality assessment module to provide a predicted average IoU score for each slice. They interactively selected slices with the lowest scores in each volume for annotation.
@@ 3.1.4 Uncertainty-aware Model
In the above sections, uncertainty is derived based on the commonly used deterministic model in DL. However, some models can inherently capture uncertainty, such as VAE or probabilistic U-Net for medical image analysis (Kohl et al., 2018). In this way, they no longer output a point estimate but instead a distribution of possible predictions, thus mitigating over-confidence. We refer to them as uncertainty-aware models in this survey. They only require a single forward pass of the deep model, thus significantly reducing computational and time costs during inference. Evidential deep learning (EDL) and mixture density networks (MDN) are often used for uncertainty-aware models in AL.
Evidential deep learning replaces the Softmax distribution with a Dirichlet distribution (Sensoy et al., 2018). The network’s output is interpreted as the parameters of a Dirichlet distribution, so the predictions followed the Dirichlet distribution. The Dirichlet distribution will be sharp if the model is confident about the predictions. Otherwise, it will be flat. Another advantage brought by EDL is that AU and EU are easy to obtain with a Dirichlet distribution. In chest X-ray classification, Balaram et al. (2022) modified the EDL-based AL to accommodate the multi-label setting. Specifically, they transformed the Dirichlet distribution in EDL into multiple Beta distributions, each corresponding to one class label. They then calculated the entropy of the Beta distributions as the AU for annotation. Chen et al. (2023a) proposed a federated AL method based on EDL for medical image analysis. Following the federated AL setting, they kept a global model across all clients and local models for each client. AUs of the global and local models and the EUs of the global models were used for sample selection. Park et al. (2023) introduced a model evidence head to scale the parameters of the Dirichlet distribution adaptively in object detection, which enhanced training stability. They first calculated the EU for each detection box. Then, the sample-level uncertainty was obtained through hierarchical uncertainty aggregation. Besides, Xie et al. (2022c) introduced EDL into active domain adaptation. Samples with high distribution and data uncertainties are selected for annotation, which are both based on EDL.
@@ 3.1.5 Adversarial-based Uncertainty
Adversarial samples help measure the sample’s distance to the decision boundary implicitly, while a higher distance indicates higher uncertainty. By attacking the deep models, adding carefully designed perturbations to original samples results in adversarial samples (Goodfellow et al., 2014b). The differences between adversarial and original samples are nearly indiscernible to the human eye. However, deep models would produce extremely confident but wrong predictions for adversarial samples. The reason is that adversarial attacks push the original samples to the other side of the decision boundary with minimal cost, resulting in visually negligible changes but significantly different predictions. From this perspective, the strength of adversarial attacks reflects the sample’s distance to the decision boundary (Heo et al., 2019). A small perturbation indicates that the sample is closer to the decision boundary and, thus, is considered more uncertain. Ducoffe and Precioso (2018) adopted the DeepFool algorithm (Moosavi-Dezfooli et al., 2016) for adversarial attacks. Samples with small adversarial perturbations are requested for labeling. Rangwani et al. (2021) attacked the deep model by maximizing the KL divergence between predictions of adversarial and original samples while the strength of perturbation is limited.
Adversarial training alternates between training feature extractors and classifiers with conflicting objectives, aiming to expose uncertain samples by increasing classifier disagreements. Yuan et al. (2021) and their extended work (Wan et al., 2023) implemented this with two classifiers on labeled and unlabeled datasets, first tuning classifiers while fixing the feature extractor to reveal more uncertain samples, then adjusting the feature extractor against fixed classifiers to minimize the gap between labeled and unlabeled samples. After several rounds, samples with the greatest disagreements are annotated.